# Embodied Intelligence for 3D Understanding: A Survey on 3D Scene Question Answering

### v1

@@ Abstract
3D Scene Question Answering (3D SQA) represents an interdisciplinary task that integrates 3D visual perception and natural language processing, empowering intelligent agents to comprehend and interact with complex 3D environments. Recent advances in large multimodal modelling have driven the creation of diverse datasets and spurred the development of instruction-tuning and zero-shot methods for 3D SQA. However, this rapid progress introduces challenges, particularly in achieving unified analysis and comparison across datasets and baselines. This paper presents the first comprehensive survey of 3D SQA, systematically reviewing datasets, methodologies, and evaluation metrics while highlighting critical challenges and future opportunities in dataset standardization, multimodal fusion, and task design.
@@ 1 Introduction
Visual Question Answering (VQA) expands the scope of traditional text-based question answering Rajpurkar et al. (2016) by incorporating visual content, enabling the interpretation of images Antol et al. (2015), charts Masry et al. (2022), and documents Ding et al. (2024) to deliver context-aware responses. This capability facilitates a broader range of applications, including medical diagnostics Wu et al. (2022), financial analysis Xue et al. (2024), and assistance in academic research. Nevertheless, the growing demand of immersive 3D environments calls for even more natural and interactive question-answering systems. 3D Scene Question Answering (3D SQA) Azuma et al. (2022); Ye et al. (2021) addresses this by bridging visual perception He et al. (2016, 2017), spatial reasoning Guo et al. (2020), and language understanding in 3D environments Linghu et al. (2024), see Figure 1. Unlike traditional 3D tasks focused on object detection Qi et al. (2019); Li et al. (2023c) or segmentation Qi et al. (2017); Zou et al. (2024); He et al. (2025), 3D SQA integrates multimodal data, e.g., visual inputs and textual queries, to enable embodied systems capable of complex reasoning Szymanska et al. (2024). By leveraging spatial relationships, object interactions, and hierarchical scene structures within dynamic 3D environments, 3D SQA advances robotics, augmented reality, and autonomous navigation Huang et al. (2023), pushing the boundaries of multimodal AI and its potential in complex, real-world scenarios.
@@
Early developments in 3D SQA were driven by manually annotated datasets like ScanQA Azuma et al. (2022) and SQA Ye et al. (2021), which aligned 3D point clouds with textual queries. Recently, programmatic generation methods, such as those used in 3DVQA Etesam et al. (2022) and MSQA Linghu et al. (2024), have enabled the creation of larger datasets with richer question types. The integration of Large Vision-Language Models (LVLMs) has further automated data annotation, leading to the development of more comprehensive datasets like LEO Huang et al. (2023) and Spartun3D Zhang et al. (2024b).
@@
Methodologies have evolved alongside datasets, transitioning from closed-set approaches to LVLM-enabled techniques. Early methods Azuma et al. (2022); Ye et al. (2021) employed custom architectures combining point cloud encoders, e.g., PointNet++ Qi et al. (2017), and text encoders, e.g., BERT Kenton and Toutanova (2019), with attention-based fusion modules. However, they were constrained by predefined answer sets. The recent LVLM-based methods employ instruction-tuning Hong et al. (2023b); Huang et al. (2023) or zero-shot technique Yin et al. (2024); Linghu et al. (2024) while adapting models like GPT-4 Achiam et al. (2023), which reduces dependence on task-specific annotations. However, these methods also face challenges in ensuring dataset quality and addressing evaluation inconsistencies.
@@
To analyse the emerging challenges in 3D SQA and facilitate their systematic handling, this paper provides the first comprehensive survey of this research direction. We focus on three fundamental aspects of this area, namely; (i) the objectives of 3D SQA, (ii) datasets needed to support these objectives, and (iii) models being developed to achieve these objectives. We review the evolution of datasets and methodologies, highlighting trends in the literature, such as the shift from manual annotation to LVLM-assisted generation, and the progression from closed-set to zero-shot methods. Additionally, we discuss challenges in multimodal alignment and evaluation standardization, offering insights into the future direction of the field. The paper outline that follows an organized structure of the existing 3D SQA literature, is provided in Figure 3 in the appendix.
Notation Definition $S$ A 3D scene representation. $S^{(p)}$ Point cloud representation of the scene: $S^{(p)}=\{(x_{i},y_{i},z_{i})\mid i=1,\dots,N\}$, where $x_{i},y_{i},z_{i}\in\mathbb{R}$ are coordinates. $S^{(m)}$ A set of multi-view images: $S^{(m)}=\{I_{1},I_{2},\dots,I_{K}\}$. $Q$ Multimodal query. $Q^{(t)}$ Textual query: $Q^{(t)}=(w_{1},w_{2},\dots,w_{L})$, where $w_{i}$ is a textual token. $Q^{(e)}$ Egocentric images in the query. $Q^{(o)}$ Object-level point clouds: $Q^{(o)}=\{(x_{j},y_{j},z_{j})\mid j=1,\dots,M\}$, $x_{j},y_{j},z_{j}\in\mathbb{R}$. $T$ A textual answer: $T=(t_{1},t_{2},\dots,t_{R})$. $B$ A set of 3D bounding boxes for objects referenced in the answer: $B=\{b_{1},b_{2},\dots,b_{M}\}$. $\mathcal{F}$ The task function mapping $(S,Q)\mapsto(T,B)$.
@@ 2 Preliminaries
The 3D SQA task involves comprehending a 3D scene $S$ and a query $Q$ to produce a textual answer $T$ and, optionally, spatial information $B$, such as bounding boxes for relevant objects. The 3D scene can be represented using modalities like point clouds $S^{(p)}$, multi-view images $S^{(m)}$, or their combinations, while the query may include textual input $Q^{(t)}$, egocentric images $Q^{(e)}$, or object-level point clouds $Q^{(o)}$. The task is formally defined as $\mathcal{F}:(S,Q)\mapsto(T,B)$, bridging multimodal reasoning and spatial understanding for comprehensive 3D scene analysis. For more details on this formulation, we refer to Table 1.
@@ 3 Datasets
ScanQA Azuma et al. (2022) SCN Dai et al. (2017) 800 41K Template $S^{(p)}$$\boldsymbol{\times}$
SQA Ye et al. (2021) SCN Dai et al. (2017) 800 6K Human $S^{(p)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
FE-3DGQA Zhao et al. (2022) SCN Dai et al. (2017) 703 20K Human $S^{(p)}$$\boldsymbol{\times}$
CLEVR3D Yan et al. (2023) 3RS Wald et al. (2019) 8,771 60K Template $S^{(p)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
3DVQA Etesam et al. (2022) SCN Dai et al. (2017) 707 500K Template $S^{(p)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
SQA3D Ma et al. (2022) SCN Dai et al. (2017) 650 33.4K Human $S^{(p)}$$\boldsymbol{\times}$
ScanScribe Zhu et al. (2023) SR Chen et al. (2020)+R3D Achlioptas et al. (2020) 2,995 56K LLM-assisted $S^{(p)}$$\boldsymbol{\times}$
3DMV-VQA Hong et al. (2023a) HM3d Ramakrishnan et al. (2021) 5K 50K Template $S^{(m)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
OpenEQA Majumdar et al. (2024) SCN, HM3d Ramakrishnan et al. (2021); Yadav et al. (2023) 180 1.6K Human $S^{(m)}$$\boldsymbol{\times}$$\boldsymbol{\times}$
Spartun3D Zhang et al. (2024b) 3RS Wald et al. (2019) - 123K LLM-assisted $\{S^{(m)},S^{(p)}\}$$\boldsymbol{\times}$
MSQA Linghu et al. (2024) SCN, 3RS, ARK Baruch et al. (2021) - 254K LLM-assisted $\{S^{(m)},S^{(p)}\}$$\boldsymbol{\times}$
LEO Huang et al. (2023) SCN+3RS Wald et al. (2019) 3K 83K LLM-assisted $\{S^{(m)},S^{(p)}\}$
M3DBench Li et al. (2023b) ScanQA Azuma et al. (2022) - 320K LLM-assisted $\{S^{(m)},S^{(p)}\}$
3D-LLM Hong et al. (2023b) Objaverse Deitke et al. (2023) - 300K LLM-assisted $\{S^{(m)},S^{(p)}\}$
LAMM Yin et al. (2024) - - 186K LLM-assisted $\{S^{(m)},S^{(p)}\}$
Table 2: Comparison of 3D SQA datasets. Source abbreviations: SCN = ScanNet, 3RS = 3RScan, HME = HM3D, ARK = ARKitScenes, SR+R3D = ScanRefer + ReferIt3D. Modality abbreviations: $S^{(p)}$ = Point Cloud, $S^{(v)}$ = Video, $S^{(m)}$ = Multi-view Images. Suited: Indicates if the dataset is an Embodied 3D SQA dataset, requiring the agent to consider its state when answering.
Importance of datasets for contemporary 3D SQA can not be overemphasized. Existing datasets vary widely in scene representation, scale, and query complexity. To provide a systematic overview of the existing datasets, this section is organized as two main parts: Dataset Structure, which explores the representation and scale of scenes and queries. and QA Pair Creation, which examines methodologies for generating question-answer pairs.
@@ 3.1.1 Scene Modalities and Scale
Synthetic 3D Datasets: The development of 3D SQA began with pseudo-3D datasets that utilized synthetic environments to simulate scene-level QA tasks. For example, EmbodiedQA Das et al. (2018) generated the dataset by selecting real scenes from the SUNCG Song et al. (2017) subset within the House3D Wu et al. (2018) simulator. These datasets were validated by human annotators to ensure quality. IQA Gordon et al. (2018) expanded this effort by introducing the IQUAD V1 dataset with 75,000 questions paired with unique scene configurations, leveraging the AI2-THOR Kolve et al. (2017) environment. MP3D-EQA Wijmans et al. (2019) and MT-EQA Yu et al. (2019) further incorporated depth maps and multi-target QA tasks, respectively, while remaining confined to synthetic SUNCG Song et al. (2017) scenes.
@@
Point Cloud Datasets: The transition to real-world 3D SQA tasks was marked by the introduction of datasets based on 3D point clouds Rusu and Cousins (2011). ScanQA Azuma et al. (2022) and SQA Ye et al. (2021) established foundational benchmarks for this direction. Both datasets were constructed using ScanNet Dai et al. (2017), with ScanQA generating 41K QA pairs across 800 scenes, and SQA providing 6K manually curated QA pairs with higher linguistic accuracy. Building on these efforts, FE-3DGQA Zhao et al. (2022) selected 703 specific scenes from ScanNet and annotated 20K QA pairs, emphasizing foundational QA tasks with dense bounding box annotations to enable spatial grounding. CLEVR3D Johnson et al. (2017) utilized functional programs and text templates to generate four times the number of questions in ScanQA, introducing a broader range of attributes and question types. Subsequently, 3DVQA Etesam et al. (2022) expanded on CLEVR3D’s framework, leveraging 3D semantic scene graphs and template-based pipelines to generate questions and answers. By selecting 707 scenes, 3DVQA produced 500K QA pairs, significantly enriching task diversity and complexity. Similarly, SQA3D Ma et al. (2022) curated 33.4K manually annotated QA pairs across 650 scenes, focusing on linking queries to agent position and orientation.
@@
Multi-View Datasets: To better align with human perception, multi-view datasets have been introduced, focusing on reasoning across different perspectives rather than relying solely on single point cloud representations. In this direction, 3DMV-VQA Hong et al. (2023a) includes 5K scenes from the HM3D dataset Ramakrishnan et al. (2021), generating 50K QA pairs. The images are rendered using the Habitat framework Ramakrishnan et al. (2021); Savva et al. (2019); Szot et al. (2021), emphasizing multi-view reasoning. On the other hand, OpenEQA Majumdar et al. (2024) not only selects scenes from HM3D but also incorporates Gibson Xia et al. (2018) and ScanNet Dai et al. (2017), ultimately choosing 180 high-quality scenes with 1.6K QA pairs. Unlike other datasets, it prioritizes quality over scale, making it a significant contribution to high-quality 3D QA benchmarks.
@@
Multimodal Datasets: Recent advances in 3D SQA datasets emphasize integrating point clouds, images, and textual data to form rich multimodal representations. These approaches aim to capture spatial, semantic, and contextual cues for more comprehensive scene understanding. A notable example is Spartun3D Zhang et al. (2024b), which selects scenes from 3RScan Wald et al. (2019) and generates 123K QA pairs focused on situational tasks. Similarly, MSQA Linghu et al. (2024) builds 254K QA pairs from multimodal datasets Dai et al. (2017); Wald et al. (2019); Baruch et al. (2021), using point clouds and object images as inputs to better align with real-world embodied intelligence scenarios.
With the popularity of LLMs, instruction tuning datasets have also emerged as an important extension of multimodal datasets, enhancing the generalization capabilities of 3D SQA models by aligning 3D data with textual descriptions. For instance, ScanScribe Zhu et al. (2023) collects RGB-D scans of indoor scenes from ScanNet and 3R-Scan, incorporating diverse object instances from Objaverse Deitke et al. (2023). It uses QA pairs from ScanQA and referential expressions from ScanRefer Chen et al. (2020) and ReferIt3D Achlioptas et al. (2020), generating 56.1K object instances from 2,995 scenes through templates and GPT-3 Brown (2020). Similarly, LEO Huang et al. (2023) constructs 83K 3D-text pairs by collecting captions at object, object-in-scene, and scene levels Luo et al. (2024); Achlioptas et al. (2020); Zhu et al. (2023); Chen et al. (2021).
@@
Along similar lines, M3DBench Li et al. (2023b) leverages multiple existing and LLMs to generate 320K instruction-response pairs, enriching multimodal 3D data for a wide range of 3D-language tasks. 3D-LLM Hong et al. (2023b) creates over 300K 3D-text pairs using assets like Objaverse, ScanNet, and HM3D, while LAMM Yin et al. (2024) employs GPT-API and self-instruction methods Wang et al. (2022) to produce 186K language-image pairs and 10K language-3D pairs.

### v2

@@
[type=editor, auid=000, bioid=1, orcid= 0009-0003-4715-703X ]
[type=editor,auid=000,bioid=1, orcid = 0000-0003-1973-6766 ]
[type=editor,auid=000,bioid=1,]
[type=editor,auid=000,bioid=1,]
[type=editor,auid=000,bioid=1,]
[type=editor,auid=000,bioid=1,]
1]organization=College of Electrical and Information Engineering,Hunan University,city=Changsha, postcode=410082, state=Hunan, country=China
2]organization=School of Computing & Information Systems ,The University of Melbourne,city=Melbourne, postcode=VIC 3053, state=VIC, country=Australia
3]organization=School of Computer Science,The University of Sydney,city=Sydney, postcode=NSW 2006, state=NSW, country=Australia
4]organization=School of Artificial Intelligence ,Anhui University,city=Hefei, postcode=230601, state=Anhui, country=China
Embodied Intelligence for 3D Understanding: A Survey on 3D Scene Question Answering
Zechuan Li lizechuan@hnu.edu.cn    Hongshan Yu yuhongshancn@hotmail.com    Yihao Ding yihao.ding@sydney.edu.au    Yan Li yali3816@uni.sydney.edu.au    Yong He h.yong@hnu.edu.cn    Naveed Akhtar naveed.akhtar1@unimelb.edu.au [ [ [ [
@@ A B S T R A C T
3D Scene Question Answering (3D SQA) represents an interdisciplinary task that integrates 3D visual perception and natural language processing, empowering intelligent agents to comprehend and interact with complex 3D environments. Recent advances in large multimodal modelling have driven the creation of diverse datasets and spurred the development of instruction-tuning and zero-shot methods for 3D SQA. However, this rapid progress introduces challenges, particularly in achieving unified analysis and comparison across datasets and baselines. In this survey, we provide the first comprehensive and systematic review of 3D SQA. We organize existing work from three perspectives: datasets, methodologies, and evaluation metrics. Beyond basic categorization, we identify shared architectural patterns across methods. Our survey further synthesizes core limitations and discusses how current trends—such as instruction tuning, multimodal alignment, and zero-shot—can shape future developments. Finally, we propose a range of promising research directions covering dataset construction, task generalization, interaction modeling, and unified evaluation protocols. This work aims to serve as a foundation for future research and foster progress toward more generalizable and intelligent 3D SQA systems.
keywords:
3D Scene Question Answering 3D Visual Perception Multi modality Large Visual Language Model Instruction-tuning Zero-shot
@@ 1 Introduction
Visual Question Answering (VQA) [109, 41] expands the scope of traditional text-based question answering [79, 10, 65] by incorporating visual content, enabling the interpretation of images [4], charts [63], and documents [22] to deliver context-aware responses. This capability facilitates a broader range of applications, including medical diagnostics [97], financial analysis [100], and assistance in academic research.
While these efforts focus primarily on 2D representations, a growing body of research in 3D visual perception has highlighted the importance of understanding spatial structures and geometric relationships in real-world environments [73, 74]. Meanwhile, the emergence of embodied intelligence has brought new demands for agents that can perceive, reason, and communicate within their 3D surroundings [108]. Applications such as household robotics [7, 53], AR/VR assistants [106], and autonomous navigation [58] require systems that not only interpret static scenes, but also support interactive multimodal understanding—comprehending user queries in context.
To meet these demands, 3D SQA(Scene Question Answering) [6, 104] has emerged as a key task that unifies spatial perception and language understanding. Unlike traditional 3D tasks that focus on object recognition and classification (e.g., 3D object detection [72, 50] or segmentation [35, 119]), 3D SQA addresses this by bridging visual perception [34, 33], spatial reasoning [28], and language understanding in 3D environments [52]. See Figure 1, 3D SQA integrates multimodal data, e.g., visual inputs and textual queries, to enable embodied systems capable of complex reasoning [88]. By leveraging spatial relationships, object interactions, and hierarchical scene structures within dynamic 3D environments, 3D SQA advances robotics, augmented reality, and autonomous navigation [40], pushing the boundaries of multimodal AI and its potential in complex, real-world scenarios.
@@
Early developments in 3D SQA were driven by manually annotated datasets like ScanQA [6] and SQA [104], which aligned 3D point clouds with textual queries. Recently, programmatic generation methods, such as those used in 3DVQA [23] and MSQA [52], have enabled the creation of larger datasets with richer question types. The integration of Large Vision-Language Models (LVLMs) has further automated data annotation, leading to the development of more comprehensive datasets like LEO [40] and Spartun3D [112].
@@
Methodologies have evolved alongside datasets, transitioning from closed-set approaches to LVLM-enabled techniques. Early methods [6, 104] employed custom architectures combining point cloud encoders, e.g., PointNet++ [74], and text encoders, e.g., BERT [44], with attention-based fusion modules. However, they were constrained by predefined answer sets. The recent LVLM-based methods employ instruction-tuning [38, 40] or zero-shot technique [105, 52] while adapting models like GPT-4 [1], which reduces dependence on task-specific annotations. However, these methods also face challenges in ensuring dataset quality and addressing evaluation inconsistencies.
@@
To analyse the emerging challenges in 3D SQA and facilitate their systematic handling, this paper provides the first comprehensive survey of this research direction. We focus on three fundamental aspects of this area, namely; (i) the objectives of 3D SQA, (ii) datasets needed to support these objectives, and (iii) models being developed to achieve these objectives. We review the evolution of datasets and methodologies, highlighting trends in the literature, such as the shift from manual annotation to LVLM-assisted generation, and the progression from closed-set to zero-shot methods. Additionally, we discuss challenges in multimodal alignment and evaluation standardization, offering insights into the future direction of the field.
To provide a clear and coherent overview of the field, we organize this survey to follow the developmental trajectory of 3D SQA itself. Since this task evolved from earlier QA and VQA paradigms, its progress has been shaped by the co-evolution of datasets, evaluation benchmarks, and modeling techniques. Accordingly, Section 2 introduces the task and its foundations, followed by Section 3 on dataset construction, Section 4 on evaluation metrics, Section 5 on representative methods, and Section 6 on open challenges. An overview of this organization is provided in Figure 2.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.2cm, s sep=5pt, calign=child edge, calign child=(n_children()+1)/2, [3D SQA , text width=3cm, fill=gray!30, for tree=fill=brown!45 [Preliminaries, for tree=fill=orange!30] [Datasets, for tree=fill=blue!40 [Dataset Structure, for tree=fill=blue!20 [Scene Modalities and Scale, for tree=fill=blue!10 [Synthetic 3D Datasets] [Point Cloud Datasets] [Multi-View Datasets] [Multimodal Datasets] ] [Query Modalities and Complexity, for tree=fill=blue!10 [Basic Text Queries] [Agent-Centric Text Queries] [Multimodal Agent-Centric Queries] [Instruction-Tuned Queries] ] ] [QA Pair Creation, for tree=fill=blue!20 [Methods for QA Pair Generation [Template-Based Generation] [Manual Annotation] [LLM-Assisted Generation] ] [Question Design in 3D SQA [Task Complexity] [Situated vs. Non-Situated] [Temporal Aspec] ] ] [Evaluating LLM-Generated 3D Datasets, for tree=fill=blue!20] ] [Evaluation Metrics, for tree=fill=green!40 [Traditional metrics] [LLM-based metrics] ] [Methodological Taxonomy, for tree=fill=violet!40 [Task-Specific Methods, for tree=fill=violet!20 [Point Cloud Methods] [Multi-view and 2D-3D Methods] [Advances in Text Encoders] ] [Pretraining-Based Methods, for tree=fill=violet!20 [Traditional Pretraining Approaches] [Instruction-Tuning Methods] ] [Zero-Shot Learning Methods, for tree=fill=violet!20 [Text-Driven Approaches] [Image-Driven Approaches] [Multimodal Alignment Approaches] ] [State-of-the-arts, for tree=fill=violet!20] ] [Challenges and Future Directions, for tree=fill=red!40 [Dataset Quality ] [Zero-Shot] [Unified Evaluation] [Dynamic and Open-World, text width=4.2cm] [Interpretable and Explainable, text width=4.2cm] [Multimodal and Collaboration, text width=4.2cm] [Incorporating Temporal Dynamics, text width=4.2cm] ] ]
Figure 2: Graphical illustration of the hierarchical structure of 3D SQA literature adopted in this work. A systematic categorization is adopted for preliminaries, datasets, evaluation metrics and methodologies.
@@
Notation Definition $S$ A 3D scene representation. $S^{(p)}$ Point cloud representation of the scene: $S^{(p)}=\{\mathbf{x}_{i}\}_{i=1}^{N},\ \mathbf{x}_{i}\in\mathbb{R}^{3}$. $S^{(m)}$ Multi-view RGB image representation: $S^{(m)}=\{I_{1},I_{2},\dots,I_{K}\}$. $Q$ Multimodal query input. $Q^{(t)}$ Textual query: $Q^{(t)}=(w_{1},w_{2},\dots,w_{L})$. $Q^{(e)}$ Egocentric image(s) as part of the query. $Q^{(o)}$ Object-level point cloud fragments: $Q^{(o)}=\{\mathbf{x}_{j}\}_{j=1}^{M},\ \mathbf{x}_{j}\in\mathbb{R}^{3}$. $T$ Textual answer: $T=(t_{1},t_{2},\dots,t_{R})$. $B^{(3D)}$ Set of 3D bounding boxes: $B^{(3D)}=\{b_{1},b_{2},\dots,b_{M}\}$. $\mathcal{F}_{3D}$ 3D SQA function mapping input to output: $\mathcal{F}_{3D}:\left(S,Q\right)\mapsto\left(T,B^{(3D)}\right)$.
@@ 2 Preliminaries
To define the 3D SQA task, we briefly review the progression of question answering paradigms across increasing levels of modality. From text-based QA to visual QA and finally to 3D scene QA, this evolution reflects a growing need for multimodal and spatially grounded understanding. 3D SQA extends prior QA settings by requiring agents to reason over spatially structured scenes—represented by point clouds or multi-view images—and respond to multimodal queries that may include text, egocentric views, or object-level inputs. This section formalizes the task setting, laying the foundation for our discussion of datasets, methods, and evaluation.
Text-based QA involves answering a textual query $Q^{(t)}$ based on a given textual passage $P$, yielding a textual response $T$:
$\mathcal{F}_{{QA}}:\left(P,Q^{{(t)}}\right)\mapsto T$ (1)
Visual QA incorporates a 2D image $I$ as context, alongside a textual question. In some cases, the output includes not only an answer $T$ but also a set of 2D bounding boxes $B^{{(2D)}}$ for visual grounding:
$\mathcal{F}_{{VQA}}:\left(I,Q^{{(t)}}\right)\mapsto\left(T,B^{{(2D)}}\right)$ (2)
3D SQA further extends this formulation by introducing a 3D scene $S$, which may consist of point clouds $S^{{(p)}}$, multi-view RGB images $S^{{(m)}}$, or both. Moreover, the query $Q$ itself may be multimodal—combining natural language $Q^{{(t)}}$, egocentric observations $Q^{{(e)}}$, or object-level 3D inputs $Q^{{(o)}}$. The system predicts a textual answer $T$, and optionally, a set of 3D bounding boxes $B^{{(3D)}}$ for spatial grounding:
$\mathcal{F}_{{3D}}:\left(S,Q\right)\mapsto\left(T,B^{{(3D)}}\right)$ (3)
This evolution from text QA to embodied 3D SQA reflects a broader shift toward situated and interactive intelligence. As tasks move from textual to visual to spatial contexts, both the input modalities and the expected outputs become increasingly multimodal and grounded.
In 3D SQA, the input consists of a 3D scene $S$ and a multimodal query $Q$. The scene $S$ may include point cloud data and multi-view images. Specifically, the point cloud is denoted as:
$S^{{(p)}}=\left\{\mathbf{x}_{i}\right\}_{i=1}^{N},\quad\mathbf{x}_{i}\in\mathbb{R}^{3},$ (4)
where each $\mathbf{x}_{i}$ represents a 3D coordinate. The multi-view RGB image input is represented as:
$S^{{(m)}}=\{I_{1},I_{2},\dots,I_{K}\}.$ (5)
The query $Q$ may also be composed of multiple modalities. A textual question is represented as a sequence of tokens:
$Q^{{(t)}}=(w_{1},w_{2},\dots,w_{L}).$ (6)
In addition, the query may include egocentric visual observations $Q^{{(e)}}$, typically corresponding to first-person views of the scene, as well as object-level point cloud fragments:
$Q^{{(o)}}=\left\{\mathbf{x}_{j}\right\}_{j=1}^{M}.$ (7)
The output of the system includes a natural language answer:
$T=(t_{1},t_{2},\dots,t_{R}),$ (8)
and, optionally, a set of 3D bounding boxes indicating spatial grounding:
$B^{{(3D)}}=\{b_{1},b_{2},\dots,b_{M}\}.$ (9)
Key notations are summarized in Table 1.
@@ 3 Datasets
Table 2: Comparison of 3D SQA datasets. Source abbreviations: SCN = ScanNet, 3RS = 3RScan, HME = HM3D, ARK = ARKitScenes, SR+R3D = ScanRefer + ReferIt3D. Modality abbreviations: $S^{(p)}$ = Point Cloud, $S^{(v)}$ = Video, $S^{(m)}$ = Multi-view Images. Suited: Indicates if the dataset is an Embodied 3D SQA dataset, requiring the agent to consider its state when answering. Grounding denotes whether the dataset includes 3D bounding box annotations for grounding answers to specific objects or regions in the scene.
@@
[…]
LEO [40] SCN+3RS [94] 3K 83K LLM-assisted $\{S^{(m)},S^{(p)}\}$
M3DBench [49] ScanQA [6] - 320K LLM-assisted $\{S^{(m)},S^{(p)}\}$
3D-LLM [38] Objaverse [20] - 300K LLM-assisted $\{S^{(m)},S^{(p)}\}$
LAMM [105] - - 186K LLM-assisted $\{S^{(m)},S^{(p)}\}$
[…]
The growing need for high-level understanding and interaction in 3D environments—such as in robotics and embodied AI—has first stimulated the construction of 3D SQA datasets. Datasets are fundamental to model development and evaluation in 3D SQA. Existing datasets vary widely in scene representation, scale, and query complexity. To provide a systematic overview, this section is organized into three parts. First, Dataset Structure examines how scenes and queries are represented, emphasizing the diversity of 3D data modalities and task requirements. Second, QA Pair Creation reviews the methodologies for generating question-answer pairs, including template-based pipelines, manual annotations, and LVLM-assisted approaches. Finally, we conclude this section with Evaluating LLM-Generated 3D Datasets, where we analyze emerging trends in dataset development and highlight key characteristics required to support the continued progress of 3D SQA research.
@@ 3.1.1 Scene Modalities and Scale
Synthetic 3D Datasets: Due to the initial lack of large-scale real-world 3D point cloud data, early research on 3D Scene Question Answering (3D SQA) relied heavily on synthetic environments to simulate scene-level QA tasks. These environments enabled fully controllable scene construction, semantic labeling, and agent simulation, providing a scalable and flexible foundation for dataset generation.
For instance, EmbodiedQA [19] constructs its dataset by selecting realistic indoor layouts from a subset of SUNCG [86] scenes rendered in the House3D [98] simulator. Questions are programmatically generated using predefined functional programs and templates [43], and the corresponding answers are obtained by executing these programs within the virtual environment.Building upon this setup, IQA [26] introduced the IQUAD V1 dataset within the AI2-THOR [45] simulator. It features 75,000 questions paired with unique scene configurations and focuses on action-conditioned QA, where agents interact with the environment (e.g., opening or picking objects) to complete a task. MP3D-EQA [96] and MT-EQA [107] further incorporated depth maps and multi-target QA tasks, respectively, while remaining confined to synthetic SUNCG [86] scenes.
@@
Point Cloud Datasets: The transition to real-world 3D SQA tasks was marked by the introduction of datasets based on 3D point clouds [81, 29]. ScanQA [6] and SQA [104] established foundational benchmarks for this direction. Both datasets were constructed using ScanNet [18], with ScanQA generating 41K QA pairs across 800 scenes, and SQA providing 6K manually curated QA pairs with higher linguistic accuracy. Building on these efforts, FE-3DGQA [115] selected 703 specific scenes from ScanNet and annotated 20K QA pairs, emphasizing foundational QA tasks with dense bounding box annotations to enable spatial grounding. CLEVR3D [43] utilized functional programs and text templates to generate four times the number of questions in ScanQA, introducing a broader range of attributes and question types. Subsequently, 3DVQA [23] expanded on CLEVR3D’s framework, leveraging 3D semantic scene graphs and template-based pipelines to generate questions and answers. By selecting 707 scenes, 3DVQA produced 500K QA pairs, significantly enriching task diversity and complexity. Similarly, SQA3D [60] marked a significant advancement in agent-centric 3D QA. It curated 33.4K manually annotated QA pairs across 650 scenes, focusing on linking queries to agent position and orientation. This dataset enabled deeper exploration of tasks that integrate agent perspectives with spatial understanding. HIS-Bench [113] is the first benchmark specifically tailored for Human-In-Scene (HIS) [5, 30] understanding. As a small yet high-quality dataset, it focuses on question answering tasks that center around human–agent interactions within 3D scenes. In contrast, View2Cap [108] emphasizes spatially grounded reasoning and constructs a large-scale corpus of 550K QA pairs, making it one of the most representative benchmarks for evaluating whether models can effectively understand 3D positional information.
@@
Multi-View Datasets: To better align with human perception, multi-view datasets have been introduced, focusing on reasoning across different perspectives rather than relying solely on single point cloud representations. In this direction, 3DMV-VQA [37] includes 5K scenes from the HM3D dataset [80], generating 50K QA pairs. The images are rendered using the Habitat framework [80, 82, 87], emphasizing multi-view reasoning. On the other hand, OpenEQA [61] not only selects scenes from HM3D but also incorporates Gibson [99] and ScanNet [18], ultimately choosing 180 high-quality scenes with 1.6K QA pairs. Unlike other datasets, it prioritizes quality over scale, making it a significant contribution to high-quality 3D SQA benchmarks.
@@
Multimodal Datasets: Recent advances in 3D SQA datasets emphasize integrating point clouds, images, and textual data to form rich multimodal representations. These approaches aim to capture spatial, semantic, and contextual cues for more comprehensive scene understanding. A notable example is Spartun3D [112], which selects scenes from 3RScan [94] and generates 123K QA pairs focused on situational tasks. Similarly, MSQA [52] builds 254K QA pairs from multimodal datasets [18, 94, 9], using point clouds and object images as inputs to better align with real-world embodied intelligence scenarios. With the popularity of LLMs, instruction tuning datasets have also emerged as an important extension of multimodal datasets, enhancing the generalization capabilities of 3D SQA models by aligning 3D data with textual descriptions. For instance, ScanScribe [118] collects RGB-D scans of indoor scenes from ScanNet and 3R-Scan, incorporating diverse object instances from Objaverse [20]. It uses QA pairs from ScanQA and referential expressions from ScanRefer [12] and ReferIt3D [2], generating 56.1K object instances from 2,995 scenes through templates and GPT-3 [11]. Similarly, LEO [40] constructs 83K 3D-text pairs by collecting captions at object, object-in-scene, and scene levels [59, 2, 118, 16]. Robospatial [85] features real-world indoor and tabletop scenes captured in the form of 3D scans and egocentric images, annotated with rich spatial information relevant to robotic tasks. The dataset includes 1 million images, 5,000 3D scans, and 3 million annotated spatial relation graphs and QA pairs, making it one of the largest and most comprehensive benchmarks available to date.
@@
Along similar lines, M3DBench [49] leverages multiple existing and LLMs to generate 320K instruction-response pairs, enriching multimodal 3D data for a wide range of 3D-language tasks. 3D-LLM [38] creates over 300K 3D-text pairs using assets like Objaverse, ScanNet, and HM3D, while LAMM [105] employs GPT-API and self-instruction methods [95] to produce 186K language-image pairs and 10K language-3D pairs.